A person can hear sound waves in the frequency range 20 Hz to 20 kHz. Find the minimum and the maximum wavelength of sound that is audible to the person. The speed of sound is 360 m s-1.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. a) for maximum wavelength n = 20 Hz. As (η∝1/λ) b) for minimum wavelength, n = 20 kHz ∴ λ = 360/ (20 × 10^3) = 18 × 10^3 m = 18 mm ⇒ * = (v/n) = 360/20 = 18 m.